OpenStudy (anonymous):

A certain radioactive isotope has a half-life of approximately 900 years. How many years would be required for a given amount of this isotope to decay to 70% of that amount?

OpenStudy (anonymous):

A=A0 e^rt

OpenStudy (anonymous):

1/2 = e^r(900)

OpenStudy (anonymous):

solve for r

OpenStudy (anonymous):

can you continue and show work

OpenStudy (anonymous):

\[e^{900r}=1/2\] 900 r ln e= ln 1/2 900 r= ln 1/2

OpenStudy (anonymous):

i have an idea

OpenStudy (anonymous):

write the formula as \[Q = Q_0(\frac{1}{2})^{\frac{t}{900}}\]

OpenStudy (anonymous):

set \[.7=(\frac{1}{2})^{\frac{t}{900}}\] and solve for t

OpenStudy (anonymous):

\[ln(.7)=\frac{t}{900}\times ln(\frac{1}{2})\]

OpenStudy (anonymous):

\[t=\frac{900ln(.7)}{ln(\frac{1}{2})}\]

OpenStudy (anonymous):

if you want to do it the first way using \[Q=Q_0e^{rt}\] then first you have to solve for r, then set the result equal to 1/2 and solve for t. it will work and you will get the same answer, but it is extra work

OpenStudy (anonymous):

ok. i see thank you

OpenStudy (anonymous):

imaram started the problem for you but you still have to solve for r, replace it in the equation, set = to 1/2 and then solve for t.